OpenStudy (anonymous):

help: find the derivative of y=(lnx)^4

OpenStudy (saifoo.khan):

It's\[\frac{4\log^3(x)}{x}\]

OpenStudy (anonymous):

how do you get that?

OpenStudy (saifoo.khan):

Derivative of: lnx = 1/x

OpenStudy (saifoo.khan):

Let me send u a photo

OpenStudy (anonymous):

okay!

OpenStudy (saifoo.khan):

It's uploading, http://ge.tt/1XHRScE/v/2?c

OpenStudy (anonymous):

okayy

OpenStudy (saifoo.khan):

u may see it now

OpenStudy (anonymous):

yeah i kind of understand

OpenStudy (saifoo.khan):

(:

OpenStudy (anonymous):

lol thank you!

OpenStudy (saifoo.khan):

Welcome. ;)